OpenStudy (anonymous):

what is another way to write (2k-1)! ?

OpenStudy (anonymous):

1*2*3*4*5.......(2k-2)*(2k-1)

OpenStudy (anonymous):

for example 5!=1*2*3*4*5=120

OpenStudy (anonymous):

for instance, (k+1)! = (k+1) * k! how does it work in the opposite direction, or does it?

OpenStudy (anonymous):

yes it does.

OpenStudy (anonymous):

so 2k! * what = (2k-1)! ?

OpenStudy (anonymous):

say you have k*k!=(k+1)*k!-k!=(k+1)!-k!

OpenStudy (anonymous):

no 2k!*1/2k =(2K-1)!

OpenStudy (anonymous):

just like the example given by dipankarstudy 5! = 5*4*3*2*1 = 5*4! = 5* (5-1)!

OpenStudy (anonymous):

it is from the basic definition of factorial.

OpenStudy (anonymous):

ah kk thank you

OpenStudy (anonymous):

dindatc my name is dipankar.

OpenStudy (anonymous):

(2k-1)! = (2k-1)(2k-2)(2k-3)(2k-4)........ (2k-1)(2k-2)(2k-3)(2k-4)........ can be written as-> (2k-1)(2k-2)!

OpenStudy (anonymous):

@dipankar ok

OpenStudy (anonymous):

yes ..thank you for making the answer more clear.

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

hi